Set A Set B
0.94 1.113
2 0.89 1.16
3 0.858 1.175
4 1.08 1.155
5 1.19 1.161
6 1.33 1.15
7 1.28 1.14
8 1.11 1.185
9 1.14

A) Calculate the expected value of the Wilcoxon Rank-Sum test Statistic E(Wx) assuming the null hypothesis of equal medians being true

B) Conduct a Wilcoxon Rank sum test on the data
Answer #1

Wilcoxon Rank - Sum test statistic is given by u - u z = - ~ N(0,1) distribution. where, u= ni+ M2 (2+1) –T u = E(u) = nim }e B) Wilcoxon Rank sum test on the given data. (Y) Set B 1.113 1.175 1.155 Sample Set A (x). 10.94 0.89 0.858 1.08 1,19 6 1.3T = 6 + 7.5 + 9+10+ 11+ 12 +13 + 14. = 82.5 0,=9 22=8 u= sint 92 (821) _T u= 9x8+ 8(90 – 82.5 2 82.5 = 72 + = 72 +36 - 82.5 T